# Condensed-space methods for nonlinear programming on GPUs

### v1

@@ 2 Primal-dual interior-point method
The interior-point method (IPM) is among the most popular algorithms to solve nonlinear programs. The basis of the algorithm is to reformulate the Karush-Kuhn-Tucker (KKT) conditions of the nonlinear program as a smooth system of nonlinear equations using a homotopy method nocedal_numerical_2006 . In a standard implementation, the resulting system is solved iteratively with a Newton method (used in conjunction with a line-search method for globalization). In Section 2.1, we give a brief description of a nonlinear program. We detail in Section 2.2 the Newton step computation within each IPM iteration.
@@ 2.1 Problem formulation and KKT conditions
$\min_{x\in\mathbb{R}^{n}}\;f(x)\quad\text{subject to}\quad\left\{\begin{aligned} &g(x)=0\;,~{}h(x)\leq 0\;,\\ &x\geq 0\;,\end{aligned}\right.$ (1)
with $f:\mathbb{R}^{n}\to\mathbb{R}$ a real-valued function encoding the objective, $g:\mathbb{R}^{n}\to\mathbb{R}^{m_{e}}$ encoding the equality constraints, and $h:\mathbb{R}^{n}\to\mathbb{R}^{m_{i}}$ encoding the inequality constraints. In addition, the variable $x$ is subject to simple bounds $x\geq 0$. In what follows, we suppose that the functions $f,g,h$ are smooth and twice differentiable.
@@
$\min_{x\in\mathbb{R}^{n},~{}s\in\mathbb{R}^{m_{i}}}\;f(x)\quad\text{subject to}\quad\left\{\begin{aligned} &g(x)=0\;,~{}h(x)+s=0\;,\\ &x\geq 0\;,~{}s\geq 0\;.\end{aligned}\right.$ (2)
In (2), the inequality constraints are encoded inside the variable bounds.
We denote by $y\in\mathbb{R}^{m_{e}}$ and $z\in\mathbb{R}^{m_{i}}$ the multipliers associated resp. to the equality constraints and the inequality constraints. Similarly, we denote by $(u,v)\in\mathbb{R}^{n+m_{i}}$ the multipliers associated respectively to the bounds $x\geq 0$ and $s\geq 0$. Using the dual variable $(y,z,u,v)$, we define the Lagrangian of (2) as
$L(x,s,y,z,u,v)=f(x)+y^{\top}g(x)+z^{\top}\big{(}h(x)+s\big{)}-u^{\top}x-v^{\top}s\;.$ (3)
@@
$\nabla f(x)+\nabla g(x)^{\top}y+\nabla h(x)^{\top}z-u=0\;,$ (4a)
@@
$0\leq x\perp u\geq 0\;,$ (4e)
$0\leq s\perp v\geq 0\;.$ (4f)
The notation $x\perp u$ is a shorthand for the complementarity condition $x_{i}u_{i}=0$ (for all $i=1\cdots,n$).
The set of active constraints at a point $x$ is denoted by
@@ 2.2 Solving the KKT conditions with the interior-point method
The interior-point method aims at finding a stationary point satisfying the KKT conditions (4). The complementarity constraints (4e)-(4f) render the KKT conditions non-smooth, complicating the solution of the whole system (4). IPM uses a homotopy continuation method to solve a simplified version of (4), parameterized by a barrier parameter $\mu>0$ (nocedal_numerical_2006, , Chapter 19). For positive $(x,s,u,v)>0$, we solve the system
$F_{\mu}(x,s,y,z,u,v)=\begin{bmatrix}\nabla f(x)+\nabla g(x)^{\top}y+\nabla h(x)^{\top}z-u\\ z-v\\ g(x)\\ h(x)+s\\ Xu-\mu e\\ Sv-\mu e\end{bmatrix}=0\;.$ (6)
We introduce in (6) the diagonal matrices $X=\operatorname{diag}(x_{1},\cdots,x_{n})$ and $S=\operatorname{diag}(s_{1},\cdots,s_{m_{i}})$, along with the vector of ones $e$. As we drive the barrier parameter $\mu$ to 00, the solution of the system $F_{\mu}(x,s,y,z,u,v)=0$ tends to the solution of the KKT conditions (4).
We note that at a fixed parameter $\mu$, the function $F_{\mu}(\cdot)$ is smooth. Hence, the system (6) can be solved iteratively using a regular Newton method. For a primal-dual iterate $w_{k}:=(x_{k},s_{k},y_{k},z_{k},u_{k},v_{k})$, the next iterate is computed as $w_{k+1}=w_{k}+\alpha_{k}d_{k}$, where $d_{k}$ is a descent direction computed by solving the linear system
@@
The step $\alpha_{k}$ is computed using a line-search algorithm, in a way that ensures that the bounded variables remain positive at the next primal-dual iterate: $(x_{k+1},s_{k+1},u_{k+1},v_{k+1})>0$. Once the iterates are sufficiently close to the central path, the IPM decreases the barrier parameter $\mu$ to find a solution closer to the original KKT conditions (4).
In IPM, the bulk of the workload is the computation of the Newton step (7), which involves assembling the Jacobian $\nabla_{w}F_{\mu}(w_{k})$ and solving the linear system to compute the descent direction $d_{k}$. By writing out all the blocks, the system in (7) expands as the $6\times 6$ unreduced KKT system:
$\begin{bmatrix}W_{k}&0&G_{k}^{\top}&H_{k}^{\top}&-I&\phantom{-}0\\ 0&0&0\phantom{{}^{\top}}&I\phantom{{}^{\top}}&\phantom{-}0&-I\\ G_{k}&0&0\phantom{{}^{\top}}&0\phantom{{}^{\top}}&\phantom{-}0&\phantom{-}0\\ H_{k}&I&0\phantom{{}^{\top}}&0\phantom{{}^{\top}}&\phantom{-}0&\phantom{-}0\\ U_{k}&0&0\phantom{{}^{\top}}&0\phantom{{}^{\top}}&\phantom{-}X_{k}&\phantom{-}0\\ 0&V_{k}&0\phantom{{}^{\top}}&0\phantom{{}^{\top}}&\phantom{-}0&\phantom{-}S_{k}\end{bmatrix}\begin{bmatrix}d_{x}\\ d_{s}\\ d_{y}\\ d_{z}\\ d_{u}\\ d_{v}\end{bmatrix}=-\begin{bmatrix}\nabla_{x}L(w_{k})\\ z_{k}-v_{k}\\ g(x_{k})\\ h(x_{k})+s_{k}\\ X_{k}u_{k}-\mu e\\ S_{k}v_{k}-\mu e\end{bmatrix}\;,$ ($K_{3}$)
where we have introduced the Hessian $W_{k}=\nabla^{2}_{xx}L(w_{k})$ and the two Jacobians $G_{k}=\nabla g(x_{k})$, $H_{k}=\nabla h(x_{k})$. In addition, we define $X_{k}$, $S_{k}$, $U_{k}$ and $V_{k}$ the diagonal matrices built respectively from the vectors $x_{k}$, $s_{k}$, $u_{k}$ and $v_{k}$. Note that ($K_{3}$) can be symmetrized by performing simple block row and column operations. In what follows, we will omit the index $k$ to alleviate the notations.
@@
It is usual to remove in ($K_{3}$) the blocks associated to the bound multipliers $(u,v)$ and solve instead the regularized $4\times 4$ symmetric system, called the augmented KKT system:
$\begin{bmatrix}W+D_{x}+\delta_{w}I&0&\phantom{-}G^{\top}&\phantom{-}H^{\top}\\ 0&D_{s}+\delta_{w}I&\phantom{-}0\phantom{{}^{\top}}&\phantom{-}I\phantom{{}^{\top}}\\ G&0&-\delta_{c}I&\phantom{-}0\phantom{{}^{\top}}\\ H&I&\phantom{-}0\phantom{{}^{\top}}&-\delta_{c}I\end{bmatrix}\begin{bmatrix}d_{x}\\ d_{s}\\ d_{y}\\ d_{z}\end{bmatrix}=-\begin{bmatrix}r_{1}\\ r_{2}\\ r_{3}\\ r_{4}\end{bmatrix}\;,$ ($K_{2}$)
with the diagonal matrices $D_{x}:=X^{-1}U$ and $D_{s}:=S^{-1}V$. The vectors forming the right-hand-sides are given respectively by $r_{1}:=\nabla f(x)+\nabla g(x)^{\top}y+\nabla h(x)^{\top}z-\mu X^{-1}e$, $r_{2}:=z-\mu S^{-1}e$, $r_{3}:=g(x)$, $r_{4}:=h(x)+s$. Once ($K_{2}$) solved, we recover the updates on bound multipliers with $d_{u}=-X^{-1}(Ud_{x}-\mu e)-u$ and $d_{v}=-S^{-1}(Vd_{s}-\mu e)-v$.
[…]
As a consequence, the system ($K_{2}$) is usually factorized using an inertia-revealing $\mathrm{LBL^{T}}$ factorization duff1983multifrontal . Krylov methods are often not competitive when solving ($K_{2}$), as the block diagonal terms $D_{x}$ and $D_{s}$ are getting increasingly ill-conditioned near the solution. Their use in IPM has been limited in linear and convex quadratic programming gondzio-2012 (when paired with a suitable preconditioner). We also refer to cao2016augmented for an efficient implementation of a preconditioned conjugate gradient on GPU, for solving the Newton step arising in an augmented Lagrangian interior-point approach.
@@
where we have introduced the condensed matrix $K:=W+D_{x}+\delta_{w}I+H^{\top}D_{H}H$ and the two diagonal matrices
@@
Compared to ($K_{3}$), the diagonal matrices $D_{x}$ and $D_{s}$ introduce an additional ill-conditioning in ($K_{2}$), amplified in the condensed form ($K_{1}$): the elements in the diagonal tend to infinity if a variable converges to its bound, and to 00 if the variable is inactive. To address the numerical error arising from such ill-conditioning, most of the implementation of IPM employs Richardson iterations on the original system ($K_{3}$) to refine the solution returned by the direct sparse linear solver (see (wachter2006implementation, , Section 3.10)).
@@ 2.3 Discussion
We have obtained three different formulations of the KKT systems appearing at each IPM iteration. The original formulation ($K_{3}$) has a better conditioning than the two alternatives ($K_{2}$) and ($K_{1}$) but has a much larger size. The second formulation ($K_{2}$) is used by default in state-of-the-art nonlinear solvers wachter2006implementation ; waltz2006interior . The system ($K_{2}$) is usually factorized using a $\mathrm{LBL^{T}}$ factorization: for sparse matrices, the Duff and Reid multifrontal algorithm duff1983multifrontal is the favored method (as implemented in the HSL linear solvers MA27 and MA57 duff2004ma57 ). The condensed KKT system ($K_{1}$) is often discarded, as its conditioning is higher than ($K_{2}$) (implying less accurate solutions). Additionally, condensation may result in increased fill-in within the condensed system ($K_{1}$(nocedal_numerical_2006, , Section 19.3, p.571). In the worst cases ($K_{1}$) itself may become fully dense if an inequality row is completely dense (fortunately, a case rarer than in the normal equations used in linear programming). Consequently, condensation methods are not commonly utilized in practical optimization settings. To the best of our knowledge, Artelys Knitro waltz2006interior is the only solver that supports computing the descent direction with ($K_{1}$).
@@ 3 Solving KKT systems on the GPU
The GPU has emerged as a new prominent computing hardware not only for graphics-related but also for general-purpose computing. GPUs employ a SIMD formalism that yields excellent throughput for parallelizing small-scale operations. However, their utility remains limited when computational algorithms require global communication. Sparse factorization algorithms, which heavily rely on numerical pivoting, pose significant challenges for implementation on GPUs. Previous research has demonstrated that GPU-based linear solvers significantly lag behind their CPU counterparts swirydowicz2021linear ; tasseff2019exploring . One emerging strategy is to utilize sparse factorization techniques that do not necessitate numerical pivoting regev2023hykkt ; shin2023accelerating by leveraging the structure of the condensed KKT system ($K_{1}$). We present two alternative methods to solve ($K_{1}$). On the one hand, HyKKT is introduced in §3.1 and uses the hybrid strategy of Golub & Greif golub2003solving ; regev2023hykkt . On the other hand, LiftedKKT shin2023accelerating uses an equality relaxation strategy and is presented in §3.2.

### v2

@@ 2 Primal-dual interior-point method
The interior-point method (IPM) is among the most popular algorithms to solve nonlinear programs. The basis of the algorithm is to reformulate the Karush-Kuhn-Tucker (KKT) conditions of the nonlinear program as a smooth system of nonlinear equations using a homotopy method nocedal_numerical_2006 . In a standard implementation, the resulting system is solved iteratively with a Newton method (used in conjunction with a line-search method for globalization).
This section provides a brief overview of the IPM algorithm and their numerical implementation. In Section 2.1, we give a description of a nonlinear programming problem formulation. We detail in Section 2.2 the Newton step computation within each IPM iteration.
@@ 2.1 Problem formulation and KKT conditions
$\min_{x\in\mathbb{R}^{n}}\;f(x)\quad\text{subject to}\quad g(x)=0\;,~h(x)\leq 0\;,$ (1)
with $f:\mathbb{R}^{n}\to\mathbb{R}$ a real-valued function encoding the objective, $g:\mathbb{R}^{n}\to\mathbb{R}^{m_{e}}$ encoding the equality constraints, and $h:\mathbb{R}^{n}\to\mathbb{R}^{m_{i}}$ encoding the inequality constraints. In what follows, we suppose that the functions $f,g,h$ are smooth and twice differentiable.
@@
$\min_{x\in\mathbb{R}^{n},~s\in\mathbb{R}^{m_{i}}}\;f(x)\quad\text{subject to}\quad\left\{\begin{aligned} &g(x)=0\;,~h(x)+s=0\;,\\ &s\geq 0\;.\end{aligned}\right$. (2)
In (2), the inequality constraints are encoded inside the variable bounds on the slack variables. We denote by $y\in\mathbb{R}^{m_{e}}$ and $z\in\mathbb{R}^{m_{i}}$ the multipliers associated resp. to the equality constraints and the inequality constraints. Similarly, we denote by $v\in\mathbb{R}^{m_{i}}$ the multipliers associated to the bounds $s\geq 0$.
Using the dual variable $(y,z,v)$, we define the Lagrangian of (2) as
$L(x,s,y,z,v)=f(x)+y^{\top}g(x)+z^{\top}\big{(}h(x)+s\big{)}-v^{\top}s\;.$ (3)
@@
$\nabla f(x)+\nabla g(x)^{\top}y+\nabla h(x)^{\top}z=0\;,$ (4a)
@@
$0\leq s\perp v\geq 0\;.$ (4e)
The notation $s\perp v$ is a shorthand for the complementarity condition $s_{i}v_{i}=0$ (for all $i=1,\cdots,n$). The set of active constraints at a point $x$ is denoted by:
@@ 2.2 Solving the KKT conditions with the interior-point method
The interior-point method aims at finding a stationary point satisfying the KKT conditions (4). The complementarity constraints (4e) render the KKT conditions non-smooth, complicating the solution of the whole system (4). IPM uses a homotopy continuation method to solve a simplified version of (4), parameterized by a barrier parameter $\mu>0$ (nocedal_numerical_2006, , Chapter 19). For positive $(x,s,v)>0$, we solve the system
$F_{\mu}(x,s,y,z,v)=\begin{bmatrix}\nabla f(x)+\nabla g(x)^{\top}y+\nabla h(x)^{\top}z\\ z-v\\ g(x)\\ h(x)+s\\ Sv-\mu e\end{bmatrix}=0\;.$ (6)
We introduce in (6) the diagonal matrices $X=\operatorname{diag}(x_{1},\cdots,x_{n})$ and $S=\operatorname{diag}(s_{1},\cdots,s_{m_{i}})$, along with the vector of ones $e$. As we drive the barrier parameter $\mu$ to 0, the solution of the system $F_{\mu}(x,s,y,z,v)=0$ tends to the solution of the KKT conditions (4).
We note that at a fixed parameter $\mu$, the function $F_{\mu}(\cdot)$ is smooth. Hence, the system (6) can be solved iteratively using a regular Newton method. For a primal-dual iterate $w_{k}:=(x_{k},s_{k},y_{k},z_{k},v_{k})$, the next iterate is computed as $w_{k+1}=w_{k}+\alpha_{k}d_{k}$, where $d_{k}$ is a descent direction computed by solving the linear system
@@
The step $\alpha_{k}$ is computed using a line-search algorithm, in a way that ensures that the bounded variables remain positive at the next primal-dual iterate: $(x_{k+1},s_{k+1},v_{k+1})>0$. Once the iterates are sufficiently close to the central path, the IPM decreases the barrier parameter $\mu$ to find a solution closer to the original KKT conditions (4).
In IPM, the bulk of the workload is the computation of the Newton step (7), which involves assembling the Jacobian $\nabla_{w}F_{\mu}(w_{k})$ of the KKT conditions and solving the resulting linear system to compute the descent direction $d_{k}$. By writing out all the blocks, the system in (7) expands as the $5\times 5$ unreduced KKT system of the following form:
$\begin{bmatrix}W_{k}&0&G_{k}^{\top}&H_{k}^{\top}&\phantom{-}0\\ 0&0&0\phantom{{}^{\top}}&I\phantom{{}^{\top}}&-I\\ G_{k}&0&0\phantom{{}^{\top}}&0\phantom{{}^{\top}}&\phantom{-}0\\ H_{k}&I&0\phantom{{}^{\top}}&0\phantom{{}^{\top}}&\phantom{-}0\\ 0&V_{k}&0\phantom{{}^{\top}}&0\phantom{{}^{\top}}&\phantom{-}S_{k}\end{bmatrix}\begin{bmatrix}d_{x}\\ d_{s}\\ d_{y}\\ d_{z}\\ d_{v}\end{bmatrix}=-\begin{bmatrix}\nabla_{x}L(w_{k})\\ z_{k}-v_{k}\\ g(x_{k})\\ h(x_{k})+s_{k}\\ S_{k}v_{k}-\mu e\end{bmatrix}\;$, ($K_{3}$)
where we have introduced the Hessian $W_{k}=\nabla^{2}_{xx}L(w_{k})$ and the two Jacobians $G_{k}=\nabla g(x_{k})$, $H_{k}=\nabla h(x_{k})$. In addition, we define $X_{k}$, $S_{k}$, $U_{k}$ and $V_{k}$ the diagonal matrices built respectively from the vectors $x_{k}$, $s_{k}$, $u_{k}$ and $v_{k}$. Note that ($K_{3}$) can be symmetrized by performing simple block row and column operations. In what follows, we will omit the iteration counter $k$ to simplify the notations.
@@
It is usual to remove in ($K_{3}$) the blocks associated to the bound multipliers $v$ and solve instead the regularized $4\times 4$ symmetric system, called the augmented KKT system:
$\begin{bmatrix}W+\delta_{w}I&0&\phantom{-}G^{\top}&\phantom{-}H^{\top}\\ 0&D_{s}+\delta_{w}I&\phantom{-}0\phantom{{}^{\top}}&\phantom{-}I\phantom{{}^{\top}}\\ G&0&-\delta_{c}I&\phantom{-}0\phantom{{}^{\top}}\\ H&I&\phantom{-}0\phantom{{}^{\top}}&-\delta_{c}I\end{bmatrix}\begin{bmatrix}d_{x}\\ d_{s}\\ d_{y}\\ d_{z}\end{bmatrix}=-\begin{bmatrix}r_{1}\\ r_{2}\\ r_{3}\\ r_{4}\end{bmatrix}\;$, ($K_{2}$)
with the diagonal matrix $D_{s}:=S^{-1}V$. The vectors forming the right-hand-sides are given respectively by $r_{1}:=\nabla f(x)+\nabla g(x)^{\top}y+\nabla h(x)^{\top}z$, $r_{2}:=z-\mu S^{-1}e$, $r_{3}:=g(x)$, $r_{4}:=h(x)+s$. Once ($K_{2}$) is solved, we recover the updates on bound multipliers with $d_{v}=-S^{-1}(Vd_{s}-\mu e)-v$.
Note that we have added additional regularization terms $\delta_{w}\geq 0$ and $\delta_{c}\geq 0$ in ($K_{2}$), to ensure the matrix is invertible. Without the regularization terms in ($K_{2}$), the augmented KKT system is non-singular if and only if the Jacobian $J=\begin{bmatrix}G\;&\;0\\ H\;&\;I\end{bmatrix}$ is full row-rank and the matrix $\begin{bmatrix}W&0\\ 0&D_{s}\end{bmatrix}$ projected onto the null-space of the Jacobian $J$ is definite benzi2005numerical . The condition is satisfied if the inertia (defined as the respective numbers of positive, negative and zero eigenvalues) of the matrix ($K_{2}$) is $(n+m_{i},m_{i}+m_{e},0)$. We use the inertia-controlling method introduced in (wachter2006implementation, , Section 3.1) to regularize the augmented matrix by adding multiple of the identity on the diagonal of ($K_{2}$) if the inertia is not equal to $(n+m_{i},m_{e}+m_{i},0)$.
As a consequence, the system ($K_{2}$) is usually factorized using an inertia-revealing $\mathrm{LBL^{T}}$ factorization duff1983multifrontal . Krylov methods are often not competitive when solving ($K_{2}$), as the block diagonal terms $D_{s}$ are getting increasingly ill-conditioned near the solution. Their use in IPM has been limited to linear and convex quadratic programming gondzio-2012 (when paired with a suitable preconditioner). We also refer to cao2016augmented for an efficient implementation of a preconditioned conjugate gradient on GPU, for solving the Newton step arising in an augmented Lagrangian interior-point approach.
@@
where we have introduced the condensed matrix $K:=W+\delta_{w}I+H^{\top}D_{H}H$ and the two diagonal matrices
@@
This indicates that one can check the inertia of $K_{1}$ to indirectly check the inertia of $K_{2}$, which, in turn, can be used to determine the regularization parameters $\delta_{w}$ and $\delta_{c}$ needed to ensure that the solution to the KKT system $d_{k}$ is in a descent direction.
@@
Compared to ($K_{3}$), the diagonal matrix $D_{s}$ introduces an additional ill-conditioning in ($K_{2}$), which is amplified in the condensed form ($K_{1}$): the elements in the diagonal tend to infinity if a variable converges to its bound and to 0 if the variable is inactive. To address the numerical errors arising from such ill-conditioning, most implementations of IPM employ Richardson iterations on the original system ($K_{3}$) to refine the solution returned by the direct sparse linear solver (see (wachter2006implementation, , Section 3.10)).
@@ 2.3 Discussion
We have obtained three different formulations of the KKT systems appearing at each IPM iteration. The original formulation ($K_{3}$) has better conditioning than the two alternatives ($K_{2}$) and ($K_{1}$), but it is of much larger size. The second formulation ($K_{2}$), which is often called augmented KKT system, is used by default in state-of-the-art nonlinear solvers wachter2006implementation ; waltz2006interior . The system ($K_{2}$) is typically factorized using a $\mathrm{LBL^{T}}$ factorization: for sparse matrices, the Duff and Reid multifrontal algorithm duff1983multifrontal is the preferred method (as implemented in the HSL linear solvers MA27 and MA57 duff2004ma57 ). The condensed KKT system ($K_{1}$) is often discarded, as its conditioning is worse than that of ($K_{2}$) (implying less accurate solutions). Additionally, condensation may result in increased fill-in within the condensed system ($K_{1}$(nocedal_numerical_2006, , Section 19.3, p.571). In the worst cases, ($K_{1}$) itself may become fully dense if an inequality row is completely dense. Consequently, condensation methods are not commonly utilized in practical optimization settings. To the best of our knowledge, Knitro waltz2006interior is the only solver that supports computing the descent direction with ($K_{1}$).
@@ 3 Solving KKT Systems on the GPU
The GPU has emerged as a prominent computing hardware not only for graphics-related tasks but also for general-purpose computing. GPUs employ a SIMD formalism that yields excellent throughput for parallelizing small-scale operations. However, their utility remains limited when computational algorithms require global communication. Sparse factorization algorithms, which heavily rely on numerical pivoting, pose significant challenges for implementation on GPUs. Previous research has demonstrated that GPU-based linear solvers significantly lag behind their CPU counterparts swirydowicz2021linear ; tasseff2019exploring . One emerging strategy is to utilize sparse factorization techniques that do not necessitate numerical pivoting regev2023hykkt ; shin2023accelerating by leveraging the structure of the condensed KKT system ($K_{1}$).
We present two alternative methods to solve ($K_{1}$). On the one hand, HyKKT is introduced in §3.1 and uses the hybrid strategy of Golub & Greif golub2003solving ; regev2023hykkt . On the other hand, LiftedKKT shin2023accelerating employs an equality relaxation strategy and is presented in §3.2. Before we introduce these methods, we make the following observation, which allows us to simplify the setting for our analysis.
Remark 1
The primal-dual regularization parameters $(\delta_{w},\delta_{c})$ are typically determined by examining the inertia of the system $K_{2}$. In particular, the dual regularization $\delta_{c}$ is activated if zero eigenvalues are detected or if the filter line-search algorithm is entering a feasibility restoration phase (wachter2006implementation, , Section 3.3). The primal regularization $\delta_{w}$ is activated if the number of positive eigenvalues is less than the desired amount of $n+m_{i}$. For a positive $\delta_{c}$, one can solve the KKT system ($K_{1}$) by solving the following equations:
$\left(K+\frac{1}{\delta_{c}}G^{\top}G\right)d_{x}=\bar{r}_{1}-\frac{1}{\delta_{c}}\bar{r}_{2}\;,$ (10)
and by setting $d_{y}=\delta_{c}^{-1}(Gd_{x}-\bar{r}_{2})$. The system in (10) is a Schur complement of the original condensed system in ($K_{1}$) and is positive definite if the system in ($K_{1}$) has the correct inertia $(n,m_{e},0)$. That is, under the application of the standard inertia correction procedure, it can always be Cholesky factorizable. Thus, when discussing the special treatment of the condensed KKT system ($K_{1}$), we assume that the dual regularization is $\delta_{c}=0$, as $\delta_{c}>0$ will only simplify the solution procedure.